use the intercepts to graph the equation 4x-12=3y
Answer by sdmmadam@yahoo.com(530)

use the intercepts to graph the equation 4x-12=3y
4x-12=3y ----(1)
That is 4x-3y = 12
Dividing by 12
4x/12 - 3y/12 = 12/12
x/3 -y/4 = 1
That is x/(3) + y/(-4) = 1 ----(2)
comparing (1) with the standard intercept form of equation to a line
given by
x/a +y/b = 1 ----(*)
we have x - intercept = 3
and y - intercept = (-4)
This means the line cuts the x-axis at A(3,0) and the y-axis at B(0,-4)
Please draw the horizontal and the vertical lines (two mutually perpendicular) call the horizontal line the x -axis
call the vertical line the y-axis
Mark the point of intersection O, the origin.
From the origin O go (on some scale say the cm scale) a distance of 3 units along the x-axis to the right of the origin and mark A=(3,0)
From the origin O go (on the same cm scale) a distance of 4 units along the
y-axis below the origin and mark B=(0,-4)
(We have OA = 3 units and OB = 4 units)
Join the points A(3,0) and B(0,-4) and produce both ways. That is the required line